Good day,

For a different reaction, Kc= 7.45×106, kf=6.05×104s−1 and kr=8.12×10−3 s−1
. Adding a catalyst increases the forward rate constant to 2.93×107 s−1
. What is the new value of the reverse reaction constant, after adding catalyst?
I think the answer is 1.28x10^13 but unsure if the units,please help.
Thanks
1
Expert's answer
2020-08-09T08:27:05-0400

A catalyst accelerates both forward and reverse reactions, but doesn't affect the equilibrium. So kf and kr change while Kc remains constant.

From a relation between constants Kc=kf/kr:

Kr=kf/Kc,

Krcat=kfcat/Kc=2.93×10^7s-1/7.45×10^6=3.93 s-1.

The units of the constant of reverse reaction is the same as of the constant of the forward reaction, while equilibrium constant is unitless.